OpenStudy (anonymous):

If the treadmill’s initial speed is 1.7 m/s, what will its final speed be?

OpenStudy (anonymous):

b. Since Dv = vf - vi vf = 1.7 m/s + 1.4 m/s = 3.1 m/s

OpenStudy (mayankdevnani):

@nitinz570 gd job,,, but Vf² = Vi² +2as Vf=sqrt(Vi² + 2as) Here we know Vi = 1.7 m/s, a = 4.8 x 10^-3 m/s² But we don't know s, the total distance traversed on the treadmill.